undeadknight26 (undeadknight26):

An expression is shown below: f(x) = -16x2 + 22x + 3 Part A: What are the x-intercepts of the graph of the f(x)? Show your work.

undeadknight26 (undeadknight26):

this is 16 and 22 right?

jigglypuff314 (jigglypuff314):

not quite? x intercepts is when y = 0 or in this case f(x) = 0 so 0 = -16x^2 + 22x + 3 solve for x

undeadknight26 (undeadknight26):

I honestly have no idea how to do that...do i combine liker terms?

jigglypuff314 (jigglypuff314):

there are not like terms because each has a different number of x's ... if that makes any sense?... you would use the quadratic equation http://2.bp.blogspot.com/_9xX2yNGJA4E/TPm214sFVjI/AAAAAAAAADI/nVLP-mPhYCM/s1600/QuadraticFormula.JPG

OpenStudy (johnweldon1993):

No you just use the quadratic equation :)

OpenStudy (johnweldon1993):

\[\large \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (johnweldon1993):

Where a,b and c are the coefficients of your equation a = -16 b = 22 c = 3

undeadknight26 (undeadknight26):

okay so... (-22) + 22^2 - 4 x (-16) x 3 /2 x -16?

OpenStudy (johnweldon1993):

\[\large \frac{-22 \pm \sqrt{22^2 - 4(-16)(3)}}{2(-16)}\]

OpenStudy (johnweldon1993):

\[\large \frac{-22 \pm \sqrt{484 - (-192)}}{-32}\]

OpenStudy (johnweldon1993):

\[\large \frac{-22 \pm \sqrt{676}}{-32}\] Then can be broken into 2 equations... \[\large \frac{-22 + \sqrt{676}}{-32}\] and \[\large \frac{-22 - \sqrt{676}}{-32}\] Then solve them

undeadknight26 (undeadknight26):

(-22 + 26)/(-32)? (-22 - 26)/(-32)?

OpenStudy (johnweldon1993):

Correct...and completely simplified...?

undeadknight26 (undeadknight26):

4/(-32) = -.125 (-48)/(-32) = 1.5? that sounds wrong...

OpenStudy (johnweldon1993):

If it sound wrong lets check it... \[\large 0 = -16x^2 + 22x + 3\] lets try -.125 first \[\large 0 = -16(-.125)^2 + 22(-.125) + 3\] \[\large 0 = -.25 - 2.75 + 3\] \[\large 0 = -3 + 3\] \[\large 0 = 0\] \(\large \color \red{\checkmark}\)

OpenStudy (johnweldon1993):

Want to check the other?

undeadknight26 (undeadknight26):

I think i should be able to do it by my self :) thanks man!

OpenStudy (johnweldon1993):

No problem bro

undeadknight26 (undeadknight26):

yo @johnweldon1993 can you help me with like 2 more parts? they easier...

OpenStudy (johnweldon1993):

Sure what do you need man

undeadknight26 (undeadknight26):

Is the vertex of the graph of f(x) going to be a maximum or minimum? What are the coordinates of the vertex? Justify your answers and show your work.

OpenStudy (johnweldon1993):

\[\large \text{Vertex} = -\frac{b}{2a}\]

undeadknight26 (undeadknight26):

so thats: 22/2(-16)

OpenStudy (johnweldon1993):

close, notice the '-' sign in front \[\large V = -\frac{22}{2(-16)}\]

undeadknight26 (undeadknight26):

ohh i see...

OpenStudy (johnweldon1993):

So that is \[\large \frac{-22}{-32} = \frac{11}{16}\] This is the 'x' coordinate of our vertex to solve for the 'y' coordinate...you plug in 11/16 into the original equation for 'x'

undeadknight26 (undeadknight26):

so now we plug 11/16 in and find the points right?

OpenStudy (johnweldon1993):

Correct...you want to find out the 'y' value of the vertex as well...

undeadknight26 (undeadknight26):

so 11/16 = .68 f(x) = -16(.68)^2 + 22(.68) + 3

OpenStudy (johnweldon1993):

Well don't round it off just yet...you do that afterwards... 11/16 = .6875 \[\large y = -16(.6875)^2 + 22(.6875) + 3\]

undeadknight26 (undeadknight26):

thats it? thanks!

undeadknight26 (undeadknight26):

would that answer this also? What are the steps you would use to graph f(x)? Justify that you can use the answers obtained in Part A and Part B to draw the graph.

OpenStudy (johnweldon1993):

Well hang on what do you get?

OpenStudy (johnweldon1993):

So we get x = .6875 y = 10.56 right? now tell me is that a minimum of maximum?

undeadknight26 (undeadknight26):

y = -16(.6875)^2 + 22(.6875) + 3 y = -16(.47) + 22 (.6875) + 3 y = -7.52 + 15.125

undeadknight26 (undeadknight26):

y = -7.52 + 15.125 + 3 y = -7.52 + 18.125 y = 10.606?

OpenStudy (johnweldon1993):

Well yeah there was some rounding error there but regardless yes... so that is your y-coordinate.....is your vertex a minimum or a maximum? *hint* it's an upside down parabola thanks to that -16 in front*

undeadknight26 (undeadknight26):

Hmm i think its a minimum...am i right?

OpenStudy (johnweldon1993):

Not quite...upside down parabola  looks to be the maximum point on that parabola right?

undeadknight26 (undeadknight26):

Oh yesh it does look that way...

OpenStudy (johnweldon1993):

And to answer that last part you asked...how would you graph f(x) Well you would mark your 2 x-intercepts...and go up to the maximum that you found....hence...your upside down parabola

undeadknight26 (undeadknight26):

Thanks man!

OpenStudy (johnweldon1993):

No prob!

undeadknight26 (undeadknight26):

I wish i could give you more than a medal n fan because you really helped me!

OpenStudy (johnweldon1993):

Don't worry about it man